Write an equation or formula that represents the general relationship indicated. Then use the given information to find the unknown value.
W varies jointly as x and y and inversely as z. If W = 10 when x = 6 and y = 5 and z = 2, find W if x = 12 and y = 6 and z = 3.

Answer by Cromlix(4381)

W = kxy/z
k = constant of variation
To find k use values given.
W = kxy/z
10 = k*6*5/2
10 = k30/2
20 = k30
k = 20/30
k = 2/3
W = 2/3*x*y/z
W = 2/3*12*6/3
W = 48/2
W = 24.
Hope this helps.
:-)